OpenStudy (anonymous):

Jasmine is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 550 customers ordered turkey, 305 customers ordered ham, and 225 customers ordered neither turkey nor ham. Based on this data, how many of the next 1,000 customers will order both turkey and ham? Show your work, and use complete sentences.

OpenStudy (anonymous):

550+305

OpenStudy (anonymous):

855 customers according to data

OpenStudy (anonymous):

i am not sure though

OpenStudy (kropot72):

The probability of ordering turkey P(T) = 0.55 The probability of ordering ham P(H) = 0.305 The probability of ordering neither turkey nor ham is \[\large P(\bar{T} \cup \bar{H})=0.225\] therefore the probability of ordering turkey or ham is \[\large P(T \cup H)=1.000-0.225=0.775\] The probability of ordering turkey and ham can be found from the equation \[\large P(T \cup H)=P(T)+P(H)-P(T \cap H)\] which can be rearranged to give \[\large P(T \cap H)=P(T)+P(H)-P(T \cap H)\ .........(1)\] Plugging values into equation (1) we get \[\large P(T\cap H)=0.55+0.305-0.775=0.08\] The number of the next 1,000 customers who will order both turkey and ham is therefore given by \[\large 1000\times0.08=80\ customers\]